OpenStudy (anonymous):

y varies jointly with x and the square of n. When y = 25, x = 3 and n = 4. Find the value of y when x = 5 and n = 10. Round to the nearest whole number. use the formula for direct variations

OpenStudy (anonymous):

y varies jointly with x and the square of n. When y = 25, x = 3 and n = 4. Find the value of y when x = 5 and n = 10. Round to the nearest w…

OpenStudy (anonymous):

the answer would be completly different

OpenStudy (anonymous):

thats a wrong statement

OpenStudy (anonymous):

Repost.

OpenStudy (anonymous):

y varies jointly with x and the square of n. When y = 25, x = 3 and n = 4. Find the value of y when x = 5 and n = 10. Round to the nearest whole number. 260 290 320 360

OpenStudy (anonymous):

the 4 numbers are the answer options

OpenStudy (lgbasallote):

y varies jointly with x and the square of n so y = kxn^2 when y = 25, x = 3 and n= 4 so 25 = k (3)(4)^2 25 = k(3)(16) 25 = k(48) k = 25/48 when x = 5 and n =10 y = (25/48)(5)(10) are you sure about your given?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!!!